Describe how you would prepare a 2.00 L of 0.100 M KrCrO4 from 1.75 M K2CrO4 stock solution
1
Expert's answer
2018-03-31T08:48:08-0400
n (KrCrO4) = 2 * 0.1 = 0.2M
V(1.75 M KrCrO4) = 0.2 /1.75 = 0.114 L